# New Perspectives in Advancing Graph Machine Learning

### v1

@@ Abstract
We introduce a novel version of the geometric scattering transform for geometric graphs containing scalar and vector node features. This new scattering transform has desirable symmetries with respect to rigid-body roto-translations (i.e., $SE(3)$-equivariance) and may be incorporated into a geometric GNN framework. We empirically show that our equivariant scattering-based GNN achieves comparable performance to other equivariant message-passing-based GNNs at a fraction of the parameter count.
@@ 1 Introduction
The field of Geometric Deep Learning (GDL) (Bronstein et al., 2017, 2021), aims to extend the success of deep learning to data sets with geometric structure such as graphs and manifolds. Crucially, most GDL methods aim to represent the data points in a manner that respects the intrinsic symmetries of the data set. That is, two data points will be represented the same way if they differ only by an uninformative deformation such as the relabeling of the vertices of a graph or a global isometry of a manifold. In this paper, we will focus on geometric graphs where the vertices lie in Euclidean space $\mathbb{R}^{d}$ and we have both scalar- and vector-valued node features. Accordingly, following the lead of equivariant GNNs (Satorras et al., 2021; Batzner et al., 2022; Han et al., 2022; Duval et al., 2023), we will aim to produce a method which has desirable symmetries with respect to rigid motions in $\mathbb{R}^{d}$.
The most prominent success of GDL has been the rise of Graph Neural Networks (GNNs) (Kipf and Welling, 2017; Veličković et al., 2018; Xu et al., 2019; Khemani et al., 2024). Most common GNNs are based on the message-passing paradigm in which each node is represented in a manner informed by its immediate neighbors, often by local averaging. However, despite their success, message-passing networks come with some well-known limitations. Because of the use of local averaging operations, the number of layers must be kept small (typically two or three) in order to prevent the oversmoothing problem. However, this introduces a new problem, underreaching, in which case the GNN is unable to capture global structure.
@@
One possible solution to the oversmoothing vs. underreaching tradeoff is provided by the geometric scattering transform (Zou and Lerman, 2019; Gama et al., 2018; Gao et al., 2019) and associated GNNs (Min et al., 2020, 2021; Tong et al., 2024)). These networks rely on diffusion wavelets to capture the multiscale geometry of the graph in a single layer and have been shown not to suffer from oversmoothing (Wenkel et al., 2022). Additionally, using these wavelets allows for sophisticated processing of the input features with comparatively few learnable parameters.
The purpose of this paper is to introduce a novel version of diffusion wavelets, and corresponding scattering networks, for geometric graphs with both scalar-valued and vector-valued node features. We will prove that our network has desirable symmetries with respect to rotations in Euclidean space and also conduct numerical experiments showing that our method is able to achieve comparable performance to other equivariant GNNs with a substantially (often at least 90%) lower parameter count.
@@ 2 Background
Throughout, we will let $G=(V,E,w)$ be a weighted, undirected, connected graph, with weighted adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$. We assume that we are given $F$ features for each node which are stored in an $n\times F$ matrix $\mathbf{X}=\mathbf{X}^{(0)}$, so that the row, $\mathbf{X}[i,:]$, contains the features associated to $v_{i}$.
2.1 GNNs and Message Passing Networks
Most common GNNs follow the message passing paradigm in which, in each layer, the representation of each vertex is updated based on its current representation and the representations of its neighbors. As a simple illustration, we consider the Graph Convolutional Network (GCN) introduced in Kipf and Welling (2017). In this network, the layerwise update rule is given by $\mathbf{X}^{(\ell+1)}=\sigma(\widehat{\mathbf{A}}\mathbf{X}^{(\ell)}\boldsymbol{\Theta}^{(\ell)})$. Here, $\widehat{\mathbf{A}}$ is an averaging type operator which aggregates information from neighboring vertices, $\boldsymbol{\Theta}^{(\ell)}$ is a learnable $F_{\ell}\times F_{\ell+1}$ weight matrix, (where $F_{\ell}$ is the number of hidden features in the $\ell$-th layer, $F_{0}=F$), and $\sigma$ is an activation function.
It is known that the matrix $\widehat{\mathbf{A}}$ acts as a smoothing operator that averages the node features within local neighborhoods, promoting similar representations for adjacent vertices, which is a useful heuristic for, e.g., node-classification tasks on homophilous citation networks. However, after each layer, the representation of each node becomes progressively smoother. Therefore, the total number of layers must be kept small, typically two or three, to avoid oversmoothing (Nt and Maehara, 2019; Qureshi et al., 2023). However, since each message-passing layer acts locally, this creates a new problem, underreaching (Lu et al., 2024), in which the network is unable to utilize long-range interactions, which may limit its effectiveness for tasks requiring the network to capture the global structure of the graph such as predicting binding affinity or total polar surface areas of a molecule (Li et al., 2024).
2.2 Diffusion Wavelets and Geometric Scattering Networks
The geometric scattering transform (Zou and Lerman, 2019; Gama et al., 2018; Gao et al., 2019) (and associated GNNs (Min et al., 2020, 2021; Tong et al., 2024)) provides an alternative to message-passing, which, as shown in Wenkel et al. (2022), allows one to circumvent the oversmoothing vs. underreaching tradeoff via the use of diffusion wavelets to extract multiscale information from the graph and the node features. Adopting the perspective of graph signal processing (GSP) (Shuman et al., 2013; Ortega et al., 2018), we will interpret columns $\mathbf{x}$ of the feature matrix $\mathbf{X}$ as signals, i.e., functions $\mathbf{x}:V\rightarrow\mathbb{R}$, where the $\mathbf{x}[i]=\mathbf{x}(v_{i})$ is the value of $\mathbf{x}$ at $v_{i}$. We consider the lazy random walk matrix $\mathbf{P}=\frac{1}{2}(\mathbf{I}+\mathbf{D}^{-1}\mathbf{A})$ [^1] Various versions of the scattering transform (Gama et al., 2018; Gao et al., 2019; Perlmutter et al., 2023) use different normalizations of the diffusion matrix. Here, we focus on the row-normalized diffusion operator. However, all of our theory may be readily adapted to other normalizations following the analysis provided in Perlmutter et al. (2023)., where $\mathbf{D}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix.
@@ 2.2 Diffusion Wavelets and Geometric Scattering Networks
We will use $\mathbf{P}$ to construct diffusion wavelets. Letting $J$ be a positive integer, we define diffusion wavelets $\mathcal{W}_{J}=\{\mathbf{\Psi}_{j}\}_{j=0}^{J}\cup\{\mathbf{\Phi}_{J}\}$ by
$\mathbf{\Psi}_{j}$ $=\mathbf{P}^{2^{j-1}}-\mathbf{P}^{2^{j}}=\mathbf{P}^{2^{j-1}}(\mathbf{I}-\mathbf{P}^{2^{j-1}}),\quad\text{for }1\leq j\leq J,$ (1)
@@
with $\mathbf{\Psi}_{0}=\mathbf{I}-\mathbf{P}$ and $\mathbf{\Phi}_{J}=\mathbf{P}^{2^{J}}$. To understand these wavelets, observe that $\boldsymbol{\Psi}_{1}\mathbf{x}[i]=(\mathbf{P}-\mathbf{P}^{2})\mathbf{x}[i]$, the $i$-th entry of $\boldsymbol{\Psi}_{1}\mathbf{x}$, can be interpreted as describing the differences in the behavior of a signal $\mathbf{x}$ with a one-hop neighborhood of $v_{i}$ to the behavior of $\mathbf{x}$ in a two-hop neighborhood. Similarly, $\boldsymbol{\Psi}_{0}\mathbf{x}=(\mathbf{I}-\mathbf{P})\mathbf{x}$ describes how the behavior of $\mathbf{x}$ at each vertex differs from at its immediate neighbors. Collectively, the filter bank $\mathcal{W}_{J}=\{\mathbf{\Psi}_{j}\}_{j=0}^{J}\cup\{\mathbf{\Phi}_{J}\}$ acts as a multi-scale feature extractor where $\boldsymbol{\Phi}_{J}$ extracts global information from $\mathbf{x}$ (at scale $2^{J}$) and each $\boldsymbol{\Psi}_{j}$ tracks changes across different scales. From the GSP perspective $\boldsymbol{\Phi}_{J}$, is interpreted as a low-pass filter and the $\boldsymbol{\Psi}_{j}$ are interpreted as band-pass filters that highlight different frequency bands.
@@
$\mathcal{U}[j]\mathbf{x}(v)=\sigma(\boldsymbol{{\Psi}}_{j}\mathbf{x}(v)),\quad\text{and}\quad\mathcal{U}[j,j^{\prime}]\mathbf{x}(v)=\mathcal{U}[j^{\prime}]\mathcal{U}[j]\mathbf{x}(v)=\sigma(\boldsymbol{{\Psi}}_{j}\sigma(\boldsymbol{{\Psi}}_{j}\mathbf{x}(v))),$ (2)
@@
for $0\leq j\leq j^{\prime}\leq J$, where $\sigma$ is an activation function. Further, if desired, $m$-th order scattering coefficients for $m\geq 3$ can be defined similarly by ${\mathcal{U}}[j_{1},j_{2},\ldots,j_{m}]\mathbf{x}(v)={\mathcal{U}}[j_{m}]{\mathcal{U}}[j_{1},\ldots,j_{m-1}]\mathbf{x}(v)$.
@@
When used for node-level tasks, one can view the collection of all $\mathcal{U}[j]\mathbf{x}$ and $\mathcal{U}[j,j^{\prime}]\mathbf{x}$ as a new set of node-features which can then be fed into a prediction head, typically an MLP. For graph-level tasks, one typically first performs a global aggregation by computing scattering moments of the form,
$\mathcal{S}[j,q]\mathbf{x}=\|\mathcal{U}[j]\mathbf{x}\|_{q}^{q}=\sum_{i=1}^{n}\left|\mathcal{U}[j]\mathbf{x}(v_{i})\right|^{q},\>\>\>\text{and}\>\>\>\mathcal{S}[j,j^{\prime},q]\mathbf{x}=\|\mathcal{U}[j,j^{\prime}]\mathbf{x}\|_{q}^{q}=\sum_{i=1}^{n}\left|\mathcal{U}[j,j^{\prime}]\mathbf{x}(v_{i})\right|^{q},$
before applying a readout prediction head. Additionally, for both graph-level and node-level tasks, works such as Min et al. (2020, 2021); Tong et al. (2024); Johnson et al. (2025a); Wenkel et al. (2024) have incorporated the scattering transform into high-performing, end-to-end differentiable GNNs.

### v2

@@ Abstract
We introduce vector diffusion wavelets (VDWs), a novel family of wavelets inspired by the vector diffusion maps algorithm that was introduced to analyze data lying in the tangent bundle of a Riemannian manifold. We show that these wavelets may be effectively incorporated into a family of geometric graph neural networks, which we refer to as VDW-GNNs. We demonstrate that such networks are effective on synthetic point cloud data, as well as on real-world data derived from wind-field measurements and neural activity data. Theoretically, we prove that these new wavelets have desirable frame theoretic properties, similar to traditional diffusion wavelets. Additionally, we prove that these wavelets have desirable symmetries with respect to rotations and translations.
@@ 1 Introduction
The field of Geometric Deep Learning (GDL) (Bronstein et al., 2017; 2021), aims to extend the success of deep learning to data sets with geometric structure such as graphs and manifolds. Crucially, most GDL methods aim to represent the data points in a manner that respects the intrinsic structure of the data set. Furthermore, they aim to represent two data points the same way if they differ only by an uninformative deformation such as the relabeling of the vertices of a graph or a global isometry of a manifold. In this paper, we focus on geometric graphs where the vertices lie in Euclidean space $\mathbb{R}^{D}$ and we have both scalar- and vector-valued node features.
In some cases, we assume that the vertices lie upon an unknown $d$-dimensional manifold $\mathcal{M}$ with $d<D$. In these cases, we will generally assume that the features take (vector) values in the tangent bundle $\mathcal{T}\mathcal{M}$, and our method will aim to leverage this intrinsic low-dimensional structure. Additionally, when appropriate, we will show that our method has desirable symmetries with respect to rigid motions in the ambient space $\mathbb{R}^{D}$ following the lead of equivariant GNNs (Satorras et al., 2021; Han et al., 2022; Duval et al., 2023).
Most common GNNs are based on the message-passing paradigm in which each node is represented in a manner informed by its immediate neighbors, often by local averaging (Kipf and Welling, 2017; Veličković et al., 2018; Xu et al., 2019; Khemani et al., 2024). This promotes similar representations for adjacent vertices, which is a useful heuristic for, e.g., node-classification tasks on homophilous citation networks. However, after each layer, the representation of each node becomes progressively smoother. Therefore, the total number of layers must be kept small, typically to two or three, to avoid oversmoothing (Nt and Maehara, 2019; Qureshi and others, 2023), where the node representations become increasingly similar thus limiting their utility for machine learning tasks. On the other hand, since each message-passing layer acts locally, this creates a new problem: underreaching (Lu et al., 2024), where the network struggles with tasks involving long-range interactions.
@@
One possible solution to the oversmoothing versus underreaching tradeoff is provided by the use of diffusion wavelets (Coifman and Maggioni, 2006) and geometric scattering transforms (Zou and Lerman, 2019; Gama et al., 2018; Gao et al., 2019) as well as associated GNNs (Min et al., 2020; 2021; Tong et al., 2024)). The use of diffusion wavelets allows such networks to capture multiscale graph geometry in a single layer while avoiding oversmoothing (Wenkel et al., 2022). In light of these considerations, we extend diffusion wavelets and associated neural networks to the setting of vector-valued features. Specifically, the primary contributions of this paper are:
  • We introduce vector diffusion wavelets (VDWs) for graphs with vector-valued node features.
  • We prove that these wavelets have similar frame bounds to traditional diffusion wavelets and that they are equivariant with respect to rotations in the ambient space.
  • We show that (VDWs) may be effectively incorporated into VDW-GNNs, and demonstrate the utility of VDW-GNNs on synthetic and real-world data sets related to wind fields and neural-activity manifolds.
@@ 2 Background
We let $G=(V,E,w)$ be a weighted, undirected, connected graph with vertices (nodes) $V=\{v_{1},\cdots,v_{n}\}$, edges $E\subseteq{V\choose 2}=\{\{v_{i},v_{j}\}:i\neq j\}$, and weighting function $w:E\rightarrow\mathbb{R}$. We let $\mathbf{A}\in\mathbb{R}^{n\times n}$ denote the weighted adjacency matrix. For most of this work, we will focus on geometric graphs, where the vertices $v_{1},\dots,v_{n}$ lie in $\mathbb{R}^{D}$, and we will denote the $i$-th coordinate of the vertex $v_{j}$ by $v_{j}[i]$.
Adopting the perspective of graph signal processing (GSP) (Shuman et al., 2013; Ortega et al., 2018), we interpret the features associated to each node as signals, i.e., functions defined on the $V$. We assume that we are given $F_{\text{scalar}}$ scalar-valued signals and $F_{\text{vec}}$ vector-valued signals, $\mathbf{x}_{i}:V\rightarrow\mathbb{R}$, $1\leq i\leq F_{\text{scalar}}$ and $\mathbf{w}_{i}:V\rightarrow\mathbb{R}^{d}$, $1\leq i\leq F_{\text{vec}}$. When convenient, will identify the signal $\mathbf{x}_{i}$ with the vector in $\mathbb{R}^{n}$ defined by $\mathbf{x}_{i}[j]=\mathbf{x}_{i}(v_{j})$. Additionally, we may organize the scalar-values signals into an $n\times F$ matrix $\mathbf{X}=\mathbf{X}^{(0)}$, whose $i$-th column is $\mathbf{x}_{i}$ so that the row $\mathbf{X}[j,:]$ contains the features associated to $v_{j}$. Similarly, we may organize the vector-valued signals into an $n\times F_{\text{vec}}\times d$ tensor $\mathbf{W}=\mathbf{W}^{(0)}$.
2.1 Diffusion Wavelets and Geometric Scattering
The geometric scattering transform (Zou and Lerman, 2019; Gama et al., 2018; Gao et al., 2019) provides an alternative to message-passing, which, as shown in Wenkel et al. (2022), allows one to circumvent the oversmoothing vs. underreaching tradeoff via the use of diffusion wavelets to extract multiscale information.
Diffusion wavelets (Coifman and Maggioni, 2006) are constructed via the diffusion operator $\mathbf{P}$ [^1] Various versions of the geometric scattering transform use different normalizations of the diffusion matrix. All of our theory may be readily adapted to other normalizations following the analysis provided in Perlmutter et al. (2023).
$\mathbf{P}=\frac{1}{2}(\mathbf{I}+\mathbf{D}^{-1}\mathbf{A}),$ (1)
@@ 2.1 Diffusion Wavelets and Geometric Scattering
where $\mathbf{D}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix. Letting $J$ be a positive integer, we define diffusion wavelets $\mathcal{W}_{J}=\{\mathbf{\Psi}_{j}\}_{j=0}^{J}\cup\{\mathbf{\Phi}_{J}\}$ by
$\mathbf{\Psi}_{j}$ $=\mathbf{P}^{2^{j-1}}-\mathbf{P}^{2^{j}}=\mathbf{P}^{2^{j-1}}(\mathbf{I}-\mathbf{P}^{2^{j-1}})$ (2)
@@
for $1\leq j\leq J$, and $\mathbf{\Psi}_{0}=\mathbf{I}-\mathbf{P}$ and $\mathbf{\Phi}_{J}=\mathbf{P}^{2^{J}}$. To understand these wavelets, observe that for a given signal $\mathbf{x}$, $\boldsymbol{\Psi}_{1}\mathbf{x}[i]=(\mathbf{P}-\mathbf{P}^{2})\mathbf{x}[i]$, the $i$-th entry of $\boldsymbol{\Psi}_{1}\mathbf{x}$, can be interpreted as describing the differences in the behavior of a signal $\mathbf{x}$ with a one-hop neighborhood of $v_{i}$ to the behavior of $\mathbf{x}$ in a two-hop neighborhood. Similarly, $\boldsymbol{\Psi}_{0}\mathbf{x}=(\mathbf{I}-\mathbf{P})\mathbf{x}$ describes how the behavior of $\mathbf{x}$ at each vertex differs from at its immediate neighbors. Collectively, the filter bank $\mathcal{W}_{J}=\{\mathbf{\Psi}_{j}\}_{j=0}^{J}\cup\{\mathbf{\Phi}_{J}\}$ acts as a multi-scale feature extractor where $\boldsymbol{\Phi}_{J}$ extracts global information from $\mathbf{x}$ (at scale $2^{J}$) and each $\boldsymbol{\Psi}_{j}$ tracks changes across different scales. From the GSP perspective $\boldsymbol{\Phi}_{J}$, is interpreted as a low-pass filter and the $\boldsymbol{\Psi}_{j}$ are interpreted as band-pass filters that highlight different frequency bands.
@@
$\mathcal{U}[j]\mathbf{x}(v)=\sigma(\boldsymbol{{\Psi}}_{j}\mathbf{x}(v)),\quad\text{and}$ (3)
$\mathcal{U}[j,j^{\prime}]\mathbf{x}(v)=\mathcal{U}[j^{\prime}]\mathcal{U}[j]\mathbf{x}(v)=\sigma(\boldsymbol{{\Psi}}_{j}\sigma(\boldsymbol{{\Psi}}_{j}\mathbf{x}(v))),$ (4)
@@
for $0\leq j\leq j^{\prime}\leq J$, where $\sigma$ is again an activation function. Further, if desired, $m$-th order scattering coefficients for $m\geq 3$ can be defined similarly by ${\mathcal{U}}[j_{1},j_{2},\ldots,j_{m}]\mathbf{x}(v)={\mathcal{U}}[j_{m}]{\mathcal{U}}[j_{1},\ldots,j_{m-1}]\mathbf{x}(v)$.
@@
When used for node-level tasks, one can view the collection of all $\mathcal{U}[j]\mathbf{x}$ and $\mathcal{U}[j,j^{\prime}]\mathbf{x}$ as a new set of node-features that can then be fed into a prediction network. For graph-level tasks, one typically first performs a global aggregation by computing moments $\mathcal{S}[j,q]\mathbf{x}=\sum_{i=1}^{n}\left|\mathcal{U}[j]\mathbf{x}(v_{i})\right|^{q}$, and $\mathcal{S}[j,j^{\prime},q]\mathbf{x}=\sum_{i=1}^{n}\left|\mathcal{U}[j,j^{\prime}]\mathbf{x}(v_{i})\right|^{q}$ before applying a final prediction head.
Additionally, we note various works have incorporated the diffusion wavelets and scattering transforms into high-performing GNNs. For instance Wenkel et al. (2024) and Johnson et al. (2025a) use learnable combinations of the node features and the wavelets filters in each layer, whereas Tong et al. (2024) replaces with dyadic wavelets with wavelets of the form $\mathbf{P}^{t_{j}}-\mathbf{P}^{t_{j+1}}$ where the $t_{j}$ are scales that are learned through a differentiable selector matrix.
2.2 The manifold hypothesis
In this section, we consider high-dimensional point clouds $\{v_{i}\}_{i=1}^{n}\subseteq\mathbb{R}^{D}$ and review techniques that aim to find and utilize a latent low-dimensional structure by constructing a graph that may be interpreted as a discrete approximation of a latent underlying data manifold. We note that high-dimensional point clouds arise in many applications with prominent examples including single-cell data analysis Ahlmann-Eltze and Huber (2025); Kröger2024 and neural data (Kaufman et al., 2016).
Unfortunately, traditional methods often suffer from the curse of dimensionality when $D$ is large. However, it is often possible to overcome these difficulties by uncovering hidden geometric structure, thus reducing the effective dimension of the data. For instance, neural population activity, though embedded in a high-dimensional neural space where each dimension corresponds to a neuron’s activity, has been shown to occupy a low-dimensional feature space reflecting the patterns of activity that the recorded network can realistically produce during behavior (Perich et al., 2025; Sadtler et al., 2014). Setting such as this leads us to the manifold hypothesis, the assumption that the data lies on an unknown (compact, Riemannian) manifold $\mathcal{M}$, i.e., a low-dimensional subset of $\mathbb{R}^{D}$, whose intrinsic dimension $d$ is much lower than the ambient dimension $D$.
A commonly used approach for data that satisfies the manifold hypothesis is to construct a graph $G=(V,E)$, whose vertices are the data points $\{v_{i}\}_{i=1}^{n}$, that is interpreted as a discretization of the underlying manifold. For instance, the diffusion maps algorithm (Coifman and Lafon, 2006) (see also Laplacian eigenmaps (Belkin and Niyogi, 2003)) interprets the diffusion matrix $\mathbf{P}$ as a discrete approximation of an underlying manifold heat kernel. It uses the first $m$ eigenvectors and eigenvalues of $\mathbf{P}$, $\{\mathbf{u}_{i}\}_{i=1}^{m}$ and $\{\omega_{i}\}_{i=1}^{m}$ to map the data points $v_{i}$ to $\mathbb{R}^{m}$ via $v_{i}\rightarrow(\omega_{1}^{t}\mathbf{u}_{1}[i],\ldots,\omega_{m}^{t}\mathbf{u}_{m}[i]),$ where $t$ is a hyper-parameter known as diffusion time.
To connect manifold learning to geometric deep learning, we note that a natural way to construct neural networks on a compact Riemannian manifold would be to define convolutions using the spectral decomposition of the Laplace-Beltrami operator $\mathcal{L}$, paralleling popular spectral graph neural networks such as Bruna et al. (2014) and Defferrard et al. (2016) that utilize the eigendecomposition of the graph Laplacian. Papers such as Perlmutter et al. (2020); Wang et al. (2024), and Wang et al. (2022b) have analyzed the theoretical stability of such networks and works such as Chew et al. (2023); Wang et al. (2022a) have introduced numerical methods, with provable statistical consistency guarantees, for implementing such networks on point clouds sampled from an unknown manifold. Most closely related to our proposed method are Perlmutter et al. (2020); Chew et al. (2024); Johnson et al. (2025a) which have introduced networks based on diffusion wavelets (see Section 2.1).
In this paper, we extend diffusion wavelets to the case where we have vector-valued signals $\mathbf{w}:V\rightarrow\mathbb{R}^{D}$. In some cases, we will assume that these vector-valued signals take values lying in the tangent bundle $\mathcal{T}\mathcal{M}$ of an unknown manifold with each $\mathbf{w}(v_{i})$ taking values in $\mathcal{T}_{v_{i}}\mathcal{M}$, the tangent space centered at $v_{i}$. In this setting, we consider the Connection Laplacian $\Delta$ which is the natural, higher-order analog of the Laplace-Beltrami Operator $\mathcal{L}=-\text{div}\circ\nabla$. (In particular, $\Delta$ is a differential operator which acts upon functions defined on $\mathcal{T}\mathcal{M}$ whereas $\mathcal{L}$ acts on functions defined on $\mathcal{M}$.) We note that Battiloro et al. (2024) used this operator, and the associated heat semigroup $e^{-t\Delta}$, to define spectral neural networks for tangent-bundle valued data.
Our method, introduced in Section 3, extends diffusion wavelets and associated networks to this setting. We define wavelets in terms of a vector diffusion matrix $\mathbf{Q}$, based on Singer and Wu (2012), which introduced a similar operator in the context of vector diffusion maps. Similar to works such as Chew et al. (2022); Johnson et al. (2025a) which view powers of the standard diffusion operator $\mathbf{P}^{t}$ as a computationally efficient proxy for the heat semigroup $e^{-t\mathcal{L}}$ associated to the Laplace-Beltrami operator, here, we view powers of $\mathbf{Q}$ as a proxy for $e^{-t\Delta}$, the heat semigroup associated to the connection Laplacian. This approximation allows us to implement the associated vector diffusion wavelets using sparse multiplications and avoid explicit eigendecompositions, thereby significantly increasing the computational efficiency and scalability our method.